A firm is selling a single product in 2 markets (San Antonio and Houston). Demand schedules for the two markets are:

QSA = 140 – 2PSA or PSA = 70 - .5 QSA

MRSA = 70 - QSA And QH = 40 - .2PH or PH = 200 – 5QH and MRH = 200 – 10QH

Total demand is: QSA + QH at equal prices, \sum Q=120-2P+36.67-.3333P=156.67-2.33P

P = 67.15 - .4286Q, and \sum MR=67.15-.8572Q

Total cost of supplying customers is TC = $650 + $20Q.

a.) How much should be sold in each market if the firm price discriminates?

b) What prices should be charged in each market if the firm price discriminates?

c.) What are the price elasticities of demand in each market if the firm price discriminates?
Answer #1

a)

PSA = 70 - 0.5QSA

MRSA = 70 - QSA

PH = 200 - 5QH

MRH = 200 - 10QH

Q = QSA + QH

TC = 650 + 20Q

MC = 20

Profit maximising condition

MRSA = MRH = MC  

MRSA =  MC  

70 - QSA = 20  

70 - 20 = QSA  

QSA = 50

firm should sell 50 units in San Antonio market

MRH = MC  

200 - 10QH = 20  

200 - 20 = 10QH

180 = 10QH

QH = 18

firm should sell 18 units in Houston market

b)

PSA = 70 - 0.5QSA

PSA = 70 - 0.5(50) = 70 - 25 = 45

So firm should charge 45 in Sen Antonio market

PH = 200 - 5QH

PH = 200 - 5(18) = 200 - 90 = 110

Therefore firm should charge 110 in Houston market

c)

PSA = 70 - 0.5QSA

QSA = 140 - 2PSA

Price elasticity of demand

Ed = (dQ/dP)(P/Q)  

dQSA /dPSA = - 2  

Ed = (-2)(45)/50 = - 90/50 = - 1.8

|Ed| =  |-1.8| = 1.8  

Thus price elasticity of demand in San Antonio market is 2

PH = 200 - 5QH

QH = 40 - 0.2PH

dQH/dPH = - 0.2  

Ed = (- 0.2)(110/18) = - 22/18 = - 1.22

|Ed| = |- 1.22 | = 1.22

Thus price elasticity of demand in Houston market is 1.22